OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y= (2/x^2)+3

OpenStudy (anonymous):

There are multiple solutions.

OpenStudy (sirm3d):

put \(g(x)=x^2+3\) \[f(g(x))=\frac{2}{x^2+3}=\frac{2}{g(x)}\] replace \(g(x)\) by \(x\), and you'll get \(f(x)\)

OpenStudy (anonymous):

y= (2/x^2)+3 i meant

OpenStudy (sirm3d):

ooh. put \(g(x)=x^2\)

OpenStudy (sirm3d):

\[f\big(g(x)\big)=\frac{2}{x^2}+3=\frac{2}{g(x)}+3\] replace \(g(x)\) by \(x\) \[f(x)=\frac{2}{x} + 3\]